# Exploring Selective Layer Fine-Tuning in Federated Learning

### v1

@@ 1 Introduction
Foundation models [1], including BERT [2], GPT [3, 4], CLIP [5, 6], LLaMA [7], and so on [8, 9], have attracted considerable attention due to their exceptional ability in handling complex tasks [10]. When it comes to practical deployments of these models in specialized fields, fine-tuning with domain-specific data becomes critical. Nevertheless, the distributed nature of data across various users and organizations presents a challenge for centralized storage and training, as it may lead to severe privacy concerns and incur additional transmission costs. Such issues have positioned federated learning (FL) [11] as a promising paradigm to fine-tune foundation models, aligning model enhancement with privacy preservation [12, 13].
@@
Generally, FL aims to learn a global model through a collaborative process where clients perform local training and upload the parameter updates to a central server for aggregation. Given that clients have limited resources [14, 15], such as computational power, communication bandwidth, and available memory, it becomes impractical for them to fine-tune the entire foundation model. Two kinds of solutions have recently emerged to tackle this challenge. The first solution employs parameter-efficient fine-tuning techniques [16, 17, 18, 19], which introduces additional modules integrated into foundation models and updates these modules with domain-specific data while keeping the parameters of the foundation model frozen. The second one is selective model fine-tuning [20, 21, 22, 23], which only selects an impactful subset of parameters for optimization to streamline the fine-tuning process under resource constraints.
@@
This study focuses on selective model fine-tuning as it is particularly well-suited to address the inherent heterogeneity in FL, i.e., the data heterogeneity and device heterogeneity [24, 25, 26]. Specifically, clients involved in FL have non-independent and identically distributed (non-IID) data distribution and different system resources, leading to the need to customize fine-tuning strategies to such discrepancies. For example, clients with limited computation resources may opt to update only a fraction of the model, while those with sufficient resources and high-quality data prefer fine-tuning a large portion of the model to enhance performance. Selective model fine-tuning enables clients to adjust the chosen part of the model to be updated based on their own capabilities, providing a flexible and advanced solution to mitigating sub-optimal issues induced by the heterogeneity in FL.
@@
The exploration of selective model fine-tuning within the context of FL, is still in its early stages. Previous studies [23, 21, 27, 28] have concentrated on designing static strategies for subnetwork selection to improve model fine-tuning performance, without adequately considering heterogeneity among clients. To fulfill this gap, in this paper, we provide a comprehensive theoretical analysis on selective model fine-tuning in FL, focusing on a general scenario where clients are allowed to choose different layers for local training and vary their choices across different training epochs, called selective layer fine-tuning. Specifically, we formulate the optimization objective of selective layer fine-tuning in FL, and provide insights on effectively determining critical layers to achieve model convergence. Building on these insights, we further propose a novel layer selection strategy that leverages local gradients and the regulation of unified selections.
@@
  • We conduct experiments to compare the proposed layer selection strategy with baseline methods on both image and text datasets. The experimental results demonstrate the superior performance of the proposed strategy in achieving better model performance, highlighting that the proposed strategy can find critical layers for fine-tuning while considering the client heterogeneity and training dynamics in FL.  [^1] The source codes are available at https://anonymous.4open.science/r/fed_selected_tune/.
@@ 2 Related Works
Various approaches have been proposed to properly select a subset of parameters for fine-tuning foundation models within centralized training, including optimizing a non-structured mask matrix [20, 21, 22, 23, 29, 30, 31, 32] and adopting layer-wise selection strategies [31, 32, 27, 33]. For example, [20] suggests updating the parameters in a stochastic manner based on the Bernoulli distribution, while [31, 32] showcase that fine-tuning the top few layers achieves competitive model performance in downstream tasks. Moreover, [27] proposes to select layers according to their gradient statistics.
Recent studies have extended the selective fine-tuning techniques to FL scenarios [34, 35, 36, 37]. Specifically, researchers [38, 28] investigate layer-wise network decomposition to achieve selective model fine-tuning on clients. However, these works fail to offer methodologies for adaptive and dynamic layer selection that take into account the heterogeneous characteristics of clients. In addition, personalized FL algorithms [39, 40] propose to train different subnetworks on clients towards better local models. Different from previous studies, we focus on providing an in-depth understanding of selective layer fine-tuning in FL, considering the heterogeneity from the perspective of client resources and local data distributions.
@@ 3 Problem Formulation
$\min_{\theta\in\mathbb{R}^{P}}f(\theta)=\sum_{i=1}^{N}\alpha_{i}f_{i}(\theta),$ (1)
@@
$\Delta^{t}=\sum_{l\in\mathcal{L}_{t}}\sum_{i\in\mathcal{S}^{t}}w_{i,l}^{t}\sum_{k=0}^{\tau-1}g_{i,l}(\theta_{i}^{t,k};\xi_{i}^{t,k}),$ (5)
@@
Inspired by previous studies [11, 43], the aggregation weights in selective layer fine-tuning are defined based on the data ratio and the masking vectors, which are formally given as:
@@
The details of the training process are summarized in Appendix A.
@@ 4 Which Layers Should Be Selected For Fine-Tuning?
  • All clients select the same layer set for fine-tuning [39, 20, 22, 30, 32], i.e., $\mathcal{L}_{i}^{t}=\mathcal{L}_{j}^{t},\forall i\neq j$;
  • Clients fix their selections across different training epochs [44, 40], i.e., $\mathcal{L}_{i}^{t_{1}}=\mathcal{L}_{i}^{t_{2}},\forall t_{1},t_{2}\in[T]$.
@@ 4.1 Theoretical Analysis
Following previous theoretical analysis in FL [41, 42, 43], we begin with some necessary assumptions.
@@
Here we first consider a case where $\tau\!=\!1$ to simplify the analysis without affecting the insights on layer selection. The detailed analysis for the generalized case, i.e., $\tau\!>\!1$, is provided in Section B.3.
Compared with the theoretical analysis for the standard FL settings [45, 43], there exist three additional challenges in selective layer fine-tuning. Firstly, since each client only updates some layers during the local training process, the aggregated gradient is no longer an unbiased estimate of the local gradient $\nabla f_{i}(\theta^{t})$, i.e.,
@@
$\mathcal{E}_{t}\leq 2\underbrace{\Bigg{[}\Bigg{\|}\sum_{l\notin\mathcal{L}_{t}}\nabla f(\theta^{t})\Bigg{\|}^{2}\Bigg{]}}_{\mathcal{E}_{t,1}}+2\underbrace{\sum_{l\in\mathcal{L}_{t}}\chi_{\mathbf{w}_{t,l}\|\mathbf{\alpha}}\kappa_{l}^{2}}_{\mathcal{E}_{t,2}},$ (14)
@@
The proofs can be found in Section B.1.
Next we aim to analyze the impact of layer selection on the convergence of the global model. Following previous studies [46, 41], we consider an algorithm to have achieved convergence if it converges to a stationary point of the global loss function, i.e., if its expected squared gradient norm $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(\theta^{t})\right\|_{2}^{2}\right]$ is zero. The following theorem and corollary show the convergence of the proposed selective layer fine-tuning framework for FL.
@@
$\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(\theta^{t})\right\|_{2}^{2}\right]\leq\frac{2}{\eta CT}\left[f(\theta^{0})-f(\theta^{*})\right]+\frac{2\gamma\eta}{C}\sigma^{2}+\frac{1}{T}\sum_{t=1}^{T}\left(\frac{1}{\gamma\eta C}+2\right)(\mathcal{E}_{t,1}+\mathcal{E}_{t,2}).$ (15)
The proofs can be found in Section B.2.
@@ 4.2 Layer Selection Strategy
Discussions on costs
The proposed layer selective algorithm requires computing and uploading the gradients of the selected layers, whose costs are much less than those of full model fine-tuning. To be more specific, when clients select $R$ layers for fine-tuning, the proposed method achieves a reduction in transmission and computational costs to $\frac{R}{L}$ and $\frac{L\tau}{\tau+L-1}$, respectively. For example, when fine-tuning a CLIP model on the CIFAR-10 dataset, the proposed layer selection algorithm demands merely 8.3% of the transmission cost and 12% of the computation cost associated with full model fine-tuning while achieving competitive performance. The detailed analysis and numerical comparisons of the computation and communication costs are provided in Appendix C.
@@ 5.1 Settings
Datasets & Models
We conduct a series of experiments on several widely-used image classification datasets, including CIFAR-10 [47] and DomainNet [48], text classification dataset, i.e., XGLUE-NC [49], and five benchmark question-answering (QA) datasets, including SCIQ [50], OpenbookQA [51], PIQA [52], ARC-Easy and ARC-Challenge [53] datasets, as summarized in Section D.1. As for the splitting of datasets, inspired by previous studies [54, 55], we consider two commonly observed data heterogeneity among clients: (i) Label skew: adopting Dirichlet distribution to allocate data samples of the CIFAR-10 dataset; (ii) Feature skew: adopting naturally domain shift on the DomainNet, XGLUE-NC, and QA datasets. Specifically, each QA dataset is equally divided into two subsets, with each client possessing one subset of samples from one of the five datasets. We adopt the CLIP model [5] for image classification tasks and the multi-lingual XLM-Roberta-Base [56] model on the XGLUE-NC dataset. In addition, we train a LLaMA-2-7B [7] model on the QA dataset.
Server & Clients
In the experiments, we set up an FL system with a central server and $N=100$ clients. In each training epoch, the server randomly selects a subset of $20$ clients, and broadcasts the up-to-date model to these clients for local training. Besides, there are $N=10$ clients in the QA task and five clients are randomly chosen for training in each epoch. The resource budgets of clients are limited, which are quantified as the maximum number of layers they can fine-tune in the local training. For example, we use $R_{i}=1$ to indicate that the resource of client $i$ cannot afford fine-tuning more than 1 layer. The resource budgets can be identical or heterogeneous among clients.
Baselines
We compare the proposed layer selection strategy with several competitive baselines, including: (i) Top [31, 32]: Clients only fine-tune the top few layers (near the output) based on their task-specific data; (ii) Bottom [27]: Clients only fine-tune the bottom few layers (near the input) based on their task-specific data, which can be beneficial for the tasks with input-shift; (iii) Both [57]: Clients fine-tune an equal proportion of both the top and bottom layers, which shows the effectiveness for large language models; (iv) SNR [58]: Clients fine-tune the layers with higher signal-to-noise ratio (SNR) values, defined as the ratio of the mean of gradient elements to their variance; (v) RGN [59, 27]: Clients fine-tune the layers with higher relative gradient norm (RGN) values, defined as the ratio of gradient norm to the parameter norm; (vi) Moreover, we consider Full model fine-tuning, i.e., training the entire model, as the performance benchmark. More implementation details can be found in Section D.2.
@@ 5.2 Comparisons
Identical resource scenario
We first consider clients with identical computational resources, i.e., clients select the same number of layers ($R_{i}=R,\forall i\in\mathcal{N}$) for fine-tuning. The experimental results are shown in Table 1. From the model performance (accuracy) on the CIFAR-10 and DomainNet datasets, we observe that the proposed strategy demonstrates notable superiority over partial layer fine-tuning baselines. Specifically, fine-tuning only one layer of the CLIP model achieves comparable performance with tuning the entire model, since the CLIP model is sufficiently powerful to extract useful features and thus requires less training on task-specific data. This also reveals that selective layer fine-tuning well meets the performance requirement within the resources of clients.
On text datasets, including XGLUE-NC and QA, the proposed layer selection strategy and RGN demonstrate similar performance, both surpassing other baseline methods (especially Top and Both) by noticeable margins. One potential explanation for this phenomenon could be that they result in similar layer selections, indicating that updating layers with higher relative gradient norms is more beneficial than other strategies, which is consistent with previous study [27]. Moving a forward step, the proposed method adopts a flexible and dynamic layer selection strategy instead of fixed strategies, which leads to competitive performance.
Heterogeneous resource scenario
Further, we conduct experiments with heterogeneous clients, i.e., clients have different local resources and thus tend to select different numbers of layers for fine-tuning. Such a heterogeneous resource scenario is more practical [24, 25] and brings additional challenges for selective layer fine-tuning. Inspired by previous studies [41, 60], the number of layers to be fine-tuned, denoted as $R_{i}$ for client $i$, is sampled from a truncated half Normal distribution within $[1,4]$.
@@
5.3 Visualization of Selected Layers

### v2

@@ 1 Introduction
Foundation models (Bommasani et al., 2021), including BERT (Devlin et al., 2019), GPT (Radford et al., 2019; Brown et al., 2020), CLIP (Radford et al., 2021; Dosovitskiy et al., 2021), LLaMA (Touvron et al., 2023), and so on (Ramesh et al., 2021; Chowdhery et al., 2023), have attracted considerable attention due to their exceptional ability in handling complex tasks (Eloundou et al., 2023). When it comes to practical deployments of these models in specialized fields, fine-tuning with domain-specific data becomes critical. Nevertheless, the distributed nature of data across various users and organizations presents a challenge for centralized storage and training, as it may lead to severe privacy concerns and incur additional transmission costs. Such issues have positioned federated learning (FL) (McMahan et al., 2017) as a promising paradigm to fine-tune foundation models, aligning model enhancement with privacy preservation (Chen et al., 2023a; Kuang et al., 2023).
@@
Generally, FL aims to learn a global model through a collaborative process where clients perform local training and upload the parameter updates to a central server for aggregation. Given that clients have limited resources (Bonawitz et al., 2019; Imteaj et al., 2022), such as computational power, communication bandwidth, and available memory, it becomes impractical for them to fine-tune the entire foundation model. Two kinds of solutions have recently emerged to tackle this challenge. The first solution employs parameter-efficient fine-tuning techniques (Houlsby et al., 2019; Gao et al., 2021; Hu et al., 2022; Li & Liang, 2021), which introduces additional modules integrated into foundation models and updates these modules with domain-specific data while keeping the parameters of the foundation model frozen. The second one is selective model fine-tuning (Lee et al., 2019a; Xu et al., 2021; Zhang et al., 2022a; Shen et al., 2021), which only selects an impactful subset of parameters for optimization to streamline the fine-tuning process under resource constraints.
@@
This study focuses on selective model fine-tuning as it is particularly well-suited to address the inherent heterogeneity in FL, i.e., the data heterogeneity and device heterogeneity (Yang et al., 2021; Chai et al., 2019; Li et al., 2022). Specifically, clients involved in FL have non-independent and identically distributed (non-IID) data distribution and different system resources, leading to the need to customize fine-tuning strategies to such discrepancies. For example, clients with limited computation resources may opt to update only a fraction of the model, while those with sufficient resources and high-quality data prefer fine-tuning a large portion of the model to enhance performance. Selective model fine-tuning enables clients to adjust the chosen part of the model to be updated based on their own capabilities, providing a flexible and advanced solution to mitigating sub-optimal issues induced by the heterogeneity in FL.
@@
The exploration of selective model fine-tuning within the context of FL, is still in its early stages. Previous studies (Shen et al., 2021; Xu et al., 2021; Lee et al., 2022; Dun et al., 2022) have concentrated on designing static strategies for subnetwork selection to improve model fine-tuning performance, without adequately considering heterogeneity among clients. To fulfill this gap, in this paper, we provide a comprehensive theoretical analysis on selective model fine-tuning in FL, focusing on a general scenario where clients are allowed to choose different layers for local training and vary their choices across different training epochs, called selective layer fine-tuning. Specifically, we formulate the optimization objective of selective layer fine-tuning in FL, and provide insights on effectively determining critical layers to achieve model convergence. Building on these insights, we further propose a novel layer selection strategy that leverages local gradients and the regulation of unified selections.
@@
  • We conduct experiments to compare the proposed layer selection strategy with baseline methods on both image and text datasets. The experimental results demonstrate the superior performance of the proposed strategy in achieving better model performance, highlighting that the proposed strategy can find critical layers for fine-tuning while considering the client heterogeneity and training dynamics in FL.
@@ 2 Related Works
Various approaches have been proposed to properly select a subset of parameters for fine-tuning foundation models within centralized training, including optimizing a non-structured mask matrix (Lee et al., 2019a; Xu et al., 2021; Zhang et al., 2022a; Shen et al., 2021; Zaken et al., 2022; Zhang et al., 2023; Kovaleva et al., 2019; Lee et al., 2019b) and adopting layer-wise selection strategies (Kovaleva et al., 2019; Lee et al., 2019b; 2022; Kaplun et al., 2023). For example, Lee et al. (2019a) suggest updating the parameters in a stochastic manner based on the Bernoulli distribution, while Kovaleva et al. (2019); Lee et al. (2019b) showcase that fine-tuning the top few layers achieves competitive model performance in downstream tasks. Moreover, Lee et al. (2022) propose to select layers according to their gradient statistics.
Recent studies have extended the selective fine-tuning techniques to FL scenarios (Nguyen et al., 2022a; Chen et al., 2022; Hilmkil et al., 2021; Zhang et al., 2022b). Specifically, researchers (Lee et al., 2023; Dun et al., 2022) investigate layer-wise network decomposition to achieve selective model fine-tuning on clients. However, these works fail to offer methodologies for adaptive and dynamic layer selection that take into account the heterogeneous characteristics of clients. In addition, personalized FL algorithms (Pillutla et al., 2022; Chen et al., 2023b) propose to train different subnetworks on clients towards better local models. Different from previous studies, we focus on providing an in-depth understanding of selective layer fine-tuning in FL, considering the heterogeneity from the perspective of client resources and local data distributions.
@@ 3 Problem Formulation
$\vspace{-5pt}\min_{\theta\in\mathbb{R}^{P}}f(\theta)=\sum_{i=1}^{N}\alpha_{i}f_{i}(\theta),$ (1)
@@
$\Delta^{t}=\sum_{l\in\mathcal{L}_{t}}\sum_{i\in\mathcal{S}^{t}}w_{i,l}^{t}\sum_{k=0}^{\tau-1}g_{i,l}(\theta_{i}^{t,k};\xi_{i}^{t,k}),\vspace{-4pt}$ (5)
@@
Inspired by previous studies (McMahan et al., 2017; Li et al., 2020), the aggregation weights in selective layer fine-tuning are defined based on the data ratio and the masking vectors, which are formally given as:
@@
The details of the training process are summarized in Algorithm 1.
Algorithm 1 Selective Layer Fine-tuning in FL
  Input: The pre-trained global model $\theta^{0}$
  for $t=0,1,\dots,T-1$ do
     Sample a set of clients $\mathcal{S}^{t}$;
     for each client $i$ in $\mathcal{S}^{t}$ do
[…]
     end for
[…]
  end for
  Return: The global model $\theta^{T}$
@@ 4 Which Layers Should Be Selected For Fine-Tuning?
  • All clients select the same layer set for fine-tuning (Pillutla et al., 2022; Lee et al., 2019a; Zhang et al., 2022a; 2023; Lee et al., 2019b), i.e., $\mathcal{L}_{i}^{t}=\mathcal{L}_{j}^{t},\forall i\neq j$;
  • Clients fix their selections across different training epochs (Arivazhagan et al., 2019; Chen et al., 2023b), i.e., $\mathcal{L}_{i}^{t_{1}}=\mathcal{L}_{i}^{t_{2}},\forall t_{1},t_{2}\in[T]$.
@@ 4.1 Theoretical Analysis
Following previous theoretical analysis in FL (Wang et al., 2020; Karimireddy et al., 2020; Li et al., 2020), we begin with some necessary assumptions.
@@
Here we first consider a case where $\tau\!=\!1$ to simplify the analysis without affecting the insights on layer selection. The detailed analysis for the generalized case, i.e., $\tau\!>\!1$, is provided in Section A.3.
Compared with the theoretical analysis for the standard FL settings (Wang et al., 2021; Li et al., 2020), there exist three additional challenges in selective layer fine-tuning. Firstly, since each client only updates some layers during the local training process, the aggregated gradient is no longer an unbiased estimate of the local gradient $\nabla f_{i}(\theta^{t})$, i.e.,
@@
$\vspace{-5pt}\mathcal{E}_{t}\leq 2\underbrace{\Bigg{[}\Bigg{\|}\sum_{l\notin\mathcal{L}_{t}}\nabla f(\theta^{t})\Bigg{\|}^{2}\Bigg{]}}_{\mathcal{E}_{t,1}}+2\underbrace{\sum_{l\in\mathcal{L}_{t}}\chi_{\mathbf{w}_{t,l}\|\mathbf{\alpha}}\kappa_{l}^{2}}_{\mathcal{E}_{t,2}},$ (14)
@@
The proofs can be found in Section A.1.
Next we aim to analyze the impact of layer selection on the convergence of the global model. Following previous studies (Bottou et al., 2018; Wang et al., 2020), we consider an algorithm to have achieved convergence if it converges to a stationary point of the global loss function, namely, if its expected squared gradient norm $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(\theta^{t})\right\|_{2}^{2}\right]$ is zero. The following theorem and corollary show the convergence of the proposed selective layer fine-tuning framework for FL.
@@
$\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(\theta^{t})\right\|_{2}^{2}\right]\leq\frac{2}{\eta CT}\left[f(\theta^{0})\!-\!f(\theta^{*})\right]+\frac{2\gamma\eta}{C}\sigma^{2}+\frac{1}{T}\!\sum_{t=1}^{T}\left(\frac{1}{\gamma\eta C}\!+\!2\right)(\mathcal{E}_{t,1}\!+\!\mathcal{E}_{t,2}),$ (15)
where $\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{P}}f(\theta)$ is the best model with the minimal loss.
The proofs can be found in Section A.2.
@@ 4.2 Layer Selection Strategy
4.3 Discussions on Computational and Communication costs
In this section, we provide discussions on the computational cost and communication costs, considering a case where each client selects $R$ layers to fine-tune a model with $L$ layers.
Computational cost   Since all methods require the same forward operations, we focus on comparing the computational cost of backward operations in different methods. For simplicity, we assume each layer requires $b$ FLOPs of backward operations. The average computational cost of the proposed layer selection method is calculated as:
$\text{Cost}_{\text{sel}}=\underbrace{b(L-1)}_{\text{Select}}+\underbrace{bR\tau}_{\text{Fine-tune}}=b(R\tau+L-1),$ (16)
$\text{Cost}_{\text{full}}=bL\tau=\frac{L\tau}{R\tau+L-1}\text{Cost}_{\text{sel}}.$ (17)
Meanwhile, the layer selection step in the proposed method introduces slightly additional costs of $\frac{L-R}{\tau L}\text{Cost}_{\text{full}}$, which can be further reduced by evaluating the model on a smaller volume of data or making the selection decision at a lower frequency.
Communication time   The communication time is determined by the transmitted bits of different methods. The proposed method only transmits the selected layers, which is much less than uploading the full model. For example, assuming that different layers have the same number of parameters, the communication cost of the proposed method is $\frac{R}{L}$ of full model fine-tuning.
To summarize, the computational cost and communication cost (time) of the proposed method are much lower than the full model fine-tuning. In the following section, the empirical observations also verify the benefits of the proposed selective layer fine-tuning method.
@@ 5.1 Settings
Datasets & Models   We conduct a series of experiments on several widely-used image classification datasets, including CIFAR-10 (Krizhevsky & Hinton, 2009) and DomainNet (Peng et al., 2019), text classification dataset, i.e., XGLUE-NC (Liang et al., 2020), and five benchmark question-answering (QA) datasets, including SCIQ (Welbl et al., 2017), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-Easy and ARC-Challenge (Bhakthavatsalam et al., 2021) datasets. More details of the adopted datasets can be found in in Section B.1.
As for the splitting of datasets, inspired by previous studies (Zhu et al., 2021; Kim et al., 2023), we consider two commonly observed data heterogeneity among clients: (i) Label skew: adopting Dirichlet distribution to allocate data samples of the CIFAR-10 dataset; (ii) Feature skew: adopting naturally domain shift on the DomainNet, XGLUE-NC, and QA datasets. Specifically, each QA dataset is equally divided into two subsets, with each client possessing one subset of samples from one of the five datasets. We adopt the CLIP model (Radford et al., 2021) for image classification tasks and the multi-lingual XLM-Roberta-Base (Conneau et al., 2019) model on the XGLUE-NC dataset. In addition, we train a LLaMA-2-7B (Touvron et al., 2023) model on the QA dataset.
Server & Clients   In the experiments, we set up an FL system with a central server and $N=100$ clients. In each training epoch, the server randomly selects a subset of $20$ clients, and broadcasts the up-to-date model to these clients for local training. Besides, there are $N=10$ clients in the QA task and five clients are randomly chosen for training in each epoch. The resource budgets of clients are limited, which are quantified as the maximum number of layers they can fine-tune in the local training. For example, we use $R_{i}=1$ to indicate that the resource of client $i$ cannot afford fine-tuning more than 1 layer. The resource budgets can be identical or heterogeneous among clients.
Baselines   We compare the proposed layer selection strategy with several competitive baselines, including: (i) Top (Kovaleva et al., 2019; Lee et al., 2019b): Clients only fine-tune the top few layers (near the output) based on their task-specific data; (ii) Bottom (Lee et al., 2022): Clients only fine-tune the bottom few layers (near the input) based on their task-specific data, which can be beneficial for the tasks with input-shift; (iii) Both (Xiao et al., 2023): Clients fine-tune an equal proportion of both the top and bottom layers, which shows the effectiveness for large language models; (iv) SNR (Mahsereci et al., 2017): Clients fine-tune the layers with higher signal-to-noise ratio (SNR) values, defined as the ratio of the mean of gradient elements to their variance; (v) RGN (Cheng et al., 2023; Lee et al., 2022): Clients fine-tune the layers with higher relative gradient norm (RGN) values, defined as the ratio of gradient norm to the parameter norm; (vi) Moreover, we consider Full model fine-tuning, i.e., training the entire model, as the performance benchmark. More implementation details can be found in Section B.2.
@@ 5.2 Comparisons
Identical resource scenario   We first consider clients with identical computational resources, i.e., clients select the same number of layers ($R_{i}=R,\forall i\in\mathcal{N}$) for fine-tuning. The experimental results are shown in Table 1. From the model performance (accuracy) on the CIFAR-10 and DomainNet datasets, we observe that the proposed strategy demonstrates notable superiority over partial layer fine-tuning baselines. Specifically, fine-tuning only one layer of the CLIP model achieves comparable performance with tuning the entire model, since the CLIP model is sufficiently powerful to extract useful features and thus requires less training on task-specific data. This also reveals that selective layer fine-tuning well meets the performance requirement within the resources of clients.
On text datasets, including XGLUE-NC and QA, the proposed layer selection strategy and RGN demonstrate similar performance, both surpassing other baseline methods (especially Top and Both) by noticeable margins. One potential explanation for this phenomenon could be that they result in similar layer selections, indicating that updating layers with higher relative gradient norms is more beneficial than other strategies, which is consistent with previous study (Lee et al., 2022). Moving a forward step, the proposed method adopts a flexible and dynamic layer selection strategy instead of fixed strategies, which leads to competitive performance.
@@
Heterogeneous resource scenario   Further, we conduct experiments with heterogeneous clients, i.e., clients have different local resources and thus tend to select different numbers of layers for fine-tuning. Such a heterogeneous resource scenario is more practical (Yang et al., 2021; Chai et al., 2019) and brings additional challenges for selective layer fine-tuning. Inspired by previous studies (Wang et al., 2020; Nguyen et al., 2022b), the number of layers to be fine-tuned, denoted as $R_{i}$ for client $i$, is sampled from a truncated half Normal distribution within $[1,4]$.
5.3 Further Discussions
Visualization of selected layers